### v1

# Search for the production of dark Higgs in the framework of Mono-Z portal at the FCC-ee simulated electron-positron collisions at $\sqrt{s}=240$ GeV

Abstract
In the present work, we study the possible production of the dark Higgs boson ($h_{D}$) candidates, which originated from a simplified-model scenario based on the Mono-Z model, in association with a neutral gauge boson (Z). This study has been performed by studying events with dimuon plus missing transverse energy produced in the simulated electron-positron collisions at the foreseen Future Circular Collider in the Electron-Positron collision mode (FCC-ee), operating at 240 GeV center of mass energy and integrated luminosity of 10.8 ab-1. In case no new physics has been discovered, we set upper limits at a 95% confidence level on the mass of the dark Higgs.
arxiv: arXiv: V4
1 Introduction
The discovery of the Higgs boson at CERN’s Large Hadron Collider (LHC) in 2012 filled a crucial gap in the Standard Model of particle physics HiggsATLAS ; HiggsCMS . Its properties continue to be analyzed, but achieving ultimate precision requires a future collider designed to produce high quantities of Higgs bosons in a clean environment. Precise measurements could reveal insights into physics beyond the Standard Model (BSM). Many theories suggest the Higgs boson could act as a gateway to new physics, particularly through exotic decays into new light particles Higgs1 ; Higgs2 ; Higgs3 ; R1 ; atlasmonoZprime . Thus, searching for these exotic decays is vital alongside precision studies at future Higgs factories Higgs4 . While such decays into promptly decaying particles have been investigated at the LHC and proposed future facilities, new physics might also appear in other decay modes monoHiggsAtlas1 ; monoHiggsAtlas2 .
Direct searches for the process H $\rightarrow$ invisible particles have been conducted by the ATLAS HiggsAtlas24 ; HiggsAtlas25 ; HiggsAtlas26 ; HiggsAtlas27 ; HiggsAtlas28 ; HiggsAtlas29 and CMS HiggsCMS30 ; HiggsCMS31 ; HiggsCMS32 ; HiggsCMS33 ; HiggsCMS34 ; HiggsCMS35 ; HiggsCMS36 collaborations. These searches utilized data collected during Run 1 (2011–2012) and Run 2 (2015–2018). The analyses targeted channels in which the Higgs boson is produced via vector boson fusion (VBF), gluon-gluon fusion (ggH), and in association with either a vector boson (VH, where V represents either a W or Z boson) or with a top quark pair (ttH). The most stringent constraint on the branching ratio B(H → inv) has been set by the CMS experiment through the VBF channel, using data from both Run 1 and Run 2. This analysis reports a 95% confidence level (CL) upper limit of 0.18 on the branching fraction to invisible particles, with an expected value of 0.10 HiggsCMS36 . In all these analyses, it was assumed that the mass of the Higgs boson decaying into invisible particles is 125 GeV.
The LEP working group for Higgs boson searches has updated its previous combined limit for the mass of a Higgs boson that decays invisibly. This study includes the data collected in a total integrated luminosity of approximately 189-209 pb-1. No evidence was found indicating the production of invisibly decaying Higgs bosons in conjunction with a Z boson decaying into hadrons, electrons, or muons. This conclusion applies to both the overall event counts and the distributions of the discriminant variables and recoil masses. Limits on the Higgs boson mass, $m_{h}$, at the 95% confidence level are established based on analyses of hadronic and leptonic Z boson decays. Consequently, the mass ranges of 60 - 112.1 GeV from the hadronic analysis and 60 - 91.3 GeV from the leptonic analysis are excluded LEP-Higgs .
A promising approach for detecting physics beyond the Standard Model (SM) at future electron-positron colliders involves analyzing changes in the dilepton mass spectrum. These changes might appear as a new peak, which is predicted by models that incorporate neutral gauge bosons, such as Z heaveyZ or Randall-Sundrum particles extradim . Alternatively, one could observe a broad distortion in the spectrum. Such distortions could indicate the presence of Contact Interactions leptoquark ; ContactInteraction or frameworks like the ADD model ADD . To support these theories, the mass spectrum needs to show an excess or deficit of events compared to the background prediction, a result primarily driven by the production of dilepton via a Drell-Yan process.
The CMS collaboration has conducted a detailed study of signatures related to Z and Contact Interaction models ZprimeandCI . Both the ATLAS and CMS collaborations have previously searched for the massive extra neutral gauge boson Z, which is predicted by Grand Unified Theory (GUT) and Supersymmetry Extra-Gauge-bosons ; gaugeboson1 ; LR-symmetry11 ; Super-symmetry12 . However, there is currently no evidence for its existence after analyzing the full RUN II period of LHC data ZprimeandCI ; zprimeATLAS . The results from the CMS experiment have excluded the existence of Z, at a 95% Confidence Level (CL), for mass values ranging from 0.6 to 5.15 TeV, while the ATLAS experiment has ruled out mass values between 0.6 and 5.1 TeV. Additionally, both ATLAS and CMS have established stringent limits on the coupling of the Z particle to Standard Model (SM) leptons, denoted as $\texttt{g}_{l}$. Based on observations of four-muon final states, the coupling constant $\texttt{g}_{l}$ is ruled out in the range of 0.004 to 0.3, depending on the mass of the Z boson R130 ; R131 .
Previous collider experiments, like the LEP-2 lep , have shed light on important insights. For Z masses exceeding the center-of-mass energy of $\sqrt{s}=209\penalty 10000\ \text{GeV}$, the LEP collaboration set a limit of $\texttt{g}_{l}\leq 0.044\times M_{Z^{\prime}}/(\text{200 GeV})$. Conversely, for $M_{Z^{\prime}}<209\penalty 10000\ \text{GeV}$, the limit on $\texttt{g}_{l}$ remains at $\texttt{g}_{l}\leq 0.044$ R37 .
If the Z does not interact with quarks, the HL-LHC and future hadron colliders will be unable to confirm its existence. In such a scenario, electron-positron colliders, like the proposed Future Circular Collider in its Electron-Positron collision mode (FCC-ee), will play a vital role. The FCC-ee is expected to begin operations at a center of mass energy of 240 GeV ($\sqrt{s}$), with plans for a future upgrade to 365 GeV FCC-ee1 ; FCC-ee2 ; FCC-ee3 .
The initial phase of the FCC-ee is designed to be a high-luminosity and high-precision electron-positron collider, functioning as a factory for electroweak processes, top quarks, and particularly, the Higgs boson FCC-ee4 . It provides controllable energy levels and minimizes QCD background noise, making it a crucial tool in this research area.
This analysis explores a light dark Higgs ($h_{D}$) which is produced at FCC-ee in addition to a light neutral gauge bosons (Z) with a mass $M_{Z^{\prime}}\leq 90$ GeV, based on the dark Higgs (DH) simplified model within the mono-Z framework R1 . We examine simulated electron-positron collisions at the FCC-ee with 240 GeV center of mass energy, focusing on dimuon events from Z decay and large missing energy linked to dark Higgs decaying to invisible particles.
This paper is organized as follows: In Section 2, we present the theoretical framework for the mono-Z portal model. Section 3 delves into the simulation techniques employed to generate both signal and standard model background samples. Moving to Section 4, we outline the selection criteria and analysis strategy. Finally, Sections 5 and 6 provide a comprehensive overview of our results and a summary of the analysis.
2 The simplified Dark Higgs model
The mono-Z model, as discussed in R1 , is one of the top recommendations from the CMS collaboration for searching for dark Higgs at the LHC DMrecommendationsByCMS . This model encompasses the production of dark Higgs or dark matter at the LHC, as well as at future electron-positron colliders, in conjunction with a new neutral gauge boson, Z.
The mediator vector boson, Z, interacts with the Standard Model particles and dark Higgs particles in the simplified Dark Higgs (DH) scenario. The Z boson generates a light dark Higgs ($h_{D}$), which then disintegrates into a pair of DM particles ($\chi\bar{\chi}$), assuming the masses of the dark Higgs and Z are equal. The mass choice is given in table 1 and the Feynman diagram is shown in figure 1.
[…]
Figure 1: Feynman diagram of the DH model, adapted from Ref. R1
The interaction term, in the Lagrangian, between the dark Higgs and Z is given by R1
$\texttt{g}_{D}M_{Z^{\prime}}h_{D}Z^{\prime\mu}Z^{\prime}_{\mu},$
where $\texttt{g}_{D}$ denotes the coupling of Z to dark Higgs, and $M_{Z^{\prime}}$ is the $Z^{\prime}$ mass.
The interaction term, in the Lagrangian, between the SM particles and Z is given by R1
$\sum_{l}\texttt{g}_{l}\bar{\psi_{l}}\gamma_{\mu}\psi_{l}Z^{\prime\mu},$
where the coupling of Z to visible leptons is represented by $\texttt{g}_{l}$.
Table 1: The assumptions of the masses of the particles produced following the DH scenario, the light dark Higgs introduced in R1 .
Scenario       Masses assumptions
Light dark Higgs       $M_{h_{D}}=$ $\begin{cases}M_{Z^{\prime}},&M_{Z^{\prime}}<125\penalty 10000\ \text{GeV}\\ 125\penalty 10000\ \text{GeV},&M_{Z^{\prime}}>125\penalty 10000\ \text{GeV}.\end{cases}$
In the DH scenario, the only permitted decay processes are as follows: ${Z}^{\prime}\rightarrow h_{D}Z^{\prime}$, $h_{D}\rightarrow\chi\chi$, and ${Z}^{\prime}\rightarrow\mu^{+}\mu^{-}$. The total decay widths of both the ${Z}^{\prime}$ and $h_{D}$ can be calculated using the masses of ${Z}^{\prime}$ and the dark Higgs, along with the relevant coupling constants. The free parameters in this scenario include the lightest dark Higgs mass $M_{h_{D}}$, the mass of the Z boson $(M_{Z^{\prime}})$, and the couplings of Z to both leptons and dark Higgs particles, $\texttt{g}_{l}$ and $\texttt{g}_{D}$, respectively.
The CMS and ATLAS detectors have extensively searched for Z bosons over the years, confirming that heavy neutral gauge bosons do not exist in the mass range of 0.2 to 5.15 TeV. Therefore, we focus on the production of light neutral gauge bosons (Z) below 90 GeV at the FCC-ee.
Due to previous restrictions from experiments like CMS, ATLAS, and LEP-2, the value of $\texttt{g}_{l}$ is approximately 0.003 for $M_{Z^{\prime}}$ between 10 and 90 GeV R37 . In contrast, $\texttt{g}_{D}$ is set to 1.0 as suggested by the LHC Dark Matter Working Group DMrecommendations , while the masses ($M_{Z^{\prime}}$, $M_{h_{D}}$) are varied.
3 Simulation of SM backgrounds and signal samples
The SM background processes yielding muon pairs in the signal region are $Z/\gamma$ $\rightarrow\mu^{+}\mu^{-}$, and $\rightarrow\tau^{+}\tau^{-}$ production, the production of top quark pairs ($\text{t}\bar{\text{t}}\rightarrow\mu^{+}\mu^{-}+2b+2\nu$), and production of diboson ($W^{+}W^{-}\rightarrow\mu^{+}\mu^{-}+2\nu$, $ZZ\rightarrow\mu^{+}\mu^{-}+2\nu$ and $ZZ\rightarrow 4\mu$).
Table 2: The simulated SM backgrounds generated from electron-positron collisions at the FCC-ee at $\sqrt{s}=240$ GeV. Their corresponding cross-section times branching ratios for each process, and the generation order are presented. Names of these MC samples and the used generators are stated as well.
Process Deacy channel Generator $\sigma\times\text{BR}\penalty 10000\ (\text{fb})$ Order
$Z/\gamma$ $\mu^{+}\mu^{-}$ Whizard 4776.0 LO
$Z/\gamma$ $\tau^{+}\tau^{-}$ Whizard 4826.0 LO
$W$ $\mu^{+}\mu^{-}+2\nu$ Whizard 200.6 LO
$Z$ $\mu^{+}\mu^{-}+2\nu$ Whizard 5.0 LO
$Z$ $4\mu$ Whizard 0.6 LO
[…]
The DH scenario signal samples and the corresponding SM background samples are privately produced. They have been generated using the WHIZARD event generator 3.1.1 MG5 . The ISR effect was included and interfaced with Pythia 6.24 for the parton shower model and hadronization R34 . For a fast detector simulation of the IDEA detector model idea , the DELPHES package delphes was used. These were generated from electron-positron collisions at the FCC-ee with a 240 GeV center of mass energy, which corresponds to the circumstances of RUN I.
For the scenario where dark Higgs is produced along with Z, we have considered the mass assumptions as summarized in table 1. Assuming $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, table 3 shows the production cross section times branching ratios at Leading Order (LO) for various mass points of Z.
Table 3: Cross-sections times branching ratios BR($Z^{\prime}\rightarrow\mu^{+}\mu^{-}$) for the DH scenario, calculated for different values of $M_{Z^{\prime}}$ in (GeV), assuming the coupling constant $\texttt{g}_{D}$ = 1.0, and $\texttt{g}_{l}$ = 0.003, at $\sqrt{s}=240$ GeV.
$M_{Z^{\prime}}=M_{h_{D}}$ 20 30 40 50 60 70 80
$\sigma\times\text{BR}\penalty 10000\ (\text{fb})$ $7.4\times 10^{-01}$ $7.1\times 10^{-01}$ $6.7\times 10^{-01}$ $6.3\times 10^{-01}$ $6.0\times 10^{-01}$ $5.8\times 10^{-01}$ $5.7\times 10^{-01}$
The Monte Carlo simulations were used to generate the SM background samples and calculate their corresponding cross-sections for this analysis. The calculations were done in leading order and can be found in table 2. Due to the tiny value of the cross section of the leptonic decay of $t\bar{t}$, it has been excluded from the analysis.
The signal samples and SM background processes were estimated from these simulations and were normalized to their respective cross sections and an integrated luminosity of 10.8 ab-1.
An ad-hoc flat 10% uncertainty is applied to cover all possible systematic effects.
4 Event selection and backgrounds reduction
4.1 Event pre-selection
The event selection process has been designed to reconstruct a final state consisting of two muons with low transverse momentum $(p_{T})$ and missing energy accounting for the invisible particles. The selection is made by applying cuts on various kinematic parameters.
Both muons must pass a preliminary selection that includes the following criteria:
- $p^{\mu}_{T}$ (GeV) >5>5,
- $|\eta^{\mu}|$ (rad) << 2.5,
- $\Sigma_{i}p^{i}_{T}/p^{\mu}_{T}<0.1$.
Here, $\Sigma_{i}p^{i}_{T}/p^{\mu}_{T}<0.1$, represents the isolation cut in DELPHES software to reject muons produced inside jets. This cut requires that the scalar $p_{T}$ sum of all muon tracks within a cone of $\Delta R=0.5$ around the muon candidate, excluding the muon candidate itself, should not exceed 10% of the $p_{T}$ of the muon. These pre-selection cuts are listed in Table 4. Each event is selected based on two opposite-charge muons.
Table 4: Summary of cut-based event selections used in the analysis.
Pre-selection Final selection
$p^{\mu}_{T}>$ 5 GeV $p^{\mu}_{T}>$ 5 GeV
$|\eta^{\mu}|<$ 2.5 rad $|\eta^{\mu}|<$ 2.5 rad
$\Sigma_{i}p^{i}_{T}/p^{\mu}_{T}<0.1$ $\Sigma_{i}p^{i}_{T}/p^{\mu}_{T}<0.1$
$|E^{\text{miss}}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}<$ 0.4
$\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}}>$ 3 rad
$\text{cos}(\text{Angle}_{3D})<-0.8$
$\Delta R(\mu^{+}\mu^{-})<1.7$
Figure 2 illustrates the distribution of the dimuon invariant mass for events that passed the pre-selection criteria outlined in table 4, with $M_{\mu^{+}\mu^{-}}<120$ GeV. In this visualization, the red histogram represents the $Z/\gamma$ background, while the cyan histogram indicates the vector boson pair background ($WW$). The green histogram corresponds to the process $ZZ\rightarrow 2\mu 2\nu$, and the yellow histogram depicts the $ZZ\rightarrow 4\mu$ events. These histograms are presented in a stacked format. While the signals of the DH scenario, which have been generated with different masses (from 30 to 80 GeV) of the Z boson and fixing the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, are represented by different colored lines, and are overlaid.
Figure 2: The measured dimuon invariant mass spectrum, after applying pre-selection summarized in table 4, for the estimated SM backgrounds and different choices of neutral gauge boson (Z) masses generated based on the DH scenario, with $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
In $e^{+}e^{-}$ collisions, the total energy and momentum of the final states are well understood, with only minor distortions from factors like initial state radiation (ISR) and beam-energy spread (BES) resulting from synchrotron radiation and beamstrahlung effects. When the decay products of the $Z^{\prime}$ boson can be clearly identified, it is possible to precisely determine the energy and momentum of the dark Higgs particle, which in turn allows us to calculate its mass. This calculation, known as the recoil mass $(M_{rec})$, is based on energy-momentum conservation and can be achieved with excellent accuracy, regardless of how the dark Higgs boson decays Mrecoil . Then, the recoil mass is calculated from the following equation,
$M_{rec}=\sqrt{s+M_{\mu^{+}\mu^{-}}-2\sqrt{s}E^{\mu^{+}\mu^{-}}}\penalty 10000\ ,$
where $\sqrt{s}$ represents the center-of-mass energy, $E^{\mu^{+}\mu^{-}}$ denotes the energy of the dimuon pair, and $M_{\mu^{+}\mu^{-}}$ refers to the invariant mass of the dimuon pair. To avoid the influence of the Z and SM Higgs boson ($ZH$) events, we conducted our analysis with a recoil mass set below 90 GeV (i.e. $M_{rec}<90$ GeV).
Figure 3: The recoil mass spectrum, for events passing the pre-selection summarized in table 4, for the estimated SM backgrounds and different choices of the dark Higss masses generated based on the DH simplified model, with the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
In Figure 3, we present the recoil mass spectrum for events that satisfy the pre-selection criteria outlined in Table 4. The histograms illustrate the estimated Standard Model backgrounds alongside various dark Higgs masses ($M_{h_{D}}=30,40,...,80$ GeV) that were generated based on the dark Higgs simplified model. This analysis utilizes coupling constants of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
Figures 2 and 3 demonstrate that the signal samples are heavily mixed with background events across the entire dimuon invariant mass and recoil mass ranges. Hence, as outlined in the following paragraph, we need to implement stricter criteria to effectively separate the signals from Standard Model backgrounds.
4.2 Event final selection and efficiencies
Figure 4: The distributions of three variables for dimuon events, where each muon passes the low $p_{T}$ muon ID discussed in the pre-selection in table 4. The four variables are $|E^{miss}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}$ 4(a), $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{miss}}$ 4(b), $\text{cos}(\text{Angle}_{3D})$ 4(c), and $\Delta R(\mu^{+}\mu^{-})$ 4(d). The model corresponds to the DH scenario with two different values of $Z^{\prime}$ ($M_{Z^{\prime}}=$ 30 and 70 GeV) and SM backgrounds. The vertical dashed lines correspond to the chosen cut value for each variable.
In addition to the pre-selection criteria, we have applied tighter cuts based on four variables:
1. We assess the relative difference between the energy of the dimuon $(E^{\mu^{+}\mu^{-}})$ and the missing energy $(E^{\text{miss}})$. This difference is selected to be less than 0.4, defined by the condition $|E^{\text{miss}}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}<$ 0.4.
2. We calculate the azimuthal angle difference $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}}$, which is the difference between the azimuthal angles of the dimuon and the missing energy $(|\phi^{\mu^{+}\mu^{-}}-\phi^{\text{miss}}|)$. This value is required to be greater than 3.0 radians.
3. We apply a criterion on the cosine of the 3D angle between the missing energy vector and the dimuon system vector to ensure they are back-to-back, requiring that $\cos(\text{Angle}_{3D})<-0.8$.
Finally, we impose a cut on the angular separation $\Delta R(\mu^{+}\mu^{-})$ between the two opposite-sign muons, which must be less than 1.7.
The graphs in Figure 4 display the distributions of specific variables for two signal presentations of the simplified model associated with the dark Higgs scenario, corresponding to $M_{Z^{\prime}}$ values of 30 GeV and 70 GeV. Additionally, the figure includes the Standard Model (SM) backgrounds. These variables are presented for HD signal samples and SM backgrounds concerning dimuon events that satisfy the pre-selection criteria specified in Table 4, with $M_{\mu^{+}\mu^{-}}<120$ GeV.
The first variable is represented as $|E^{\text{miss}}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}$, with its graph displayed in Plot 4(a). The second variable is denoted as $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}}$, and its corresponding graph is shown in Plot 4(b). The third variable measures the variable, $\text{cos}(\text{Angle}_{3D})$, is illustrated in Plot 4(c). Lastly, the angular distance between the two muons, referred to as $\Delta R(\mu^{+}\mu^{-})$, is presented in Plot 4(d). The vertical dashed lines in these figures indicate the selected cut value for each variable.
Figure 5: Distributions of the N-1 efficiencies plotted against the transverse momentum of the leading reconstructed muon ($p^{\mu}_{T}$) for the following cuts; $|E^{miss}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}<0.4$ 5(a), $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}}>3.0$ 5(b), $\Delta R(\mu^{+}\mu^{-})<1.7$ 5(c), and $\text{cos}(\text{Angle}_{3D})<-0.8$ 5(d) for the signal in the DH scenario with $M_{Z^{\prime}}=30$ GeV and for the SM backgrounds.
The performance metrics for fine-tuning these rigorous cuts are illustrated by plotting the N-1 efficiency for each of the four criteria detailed in Table 4. To calculate the N-1 efficiency, we take the number of events that successfully pass the final selection and divide it by the number of events that would have cleared the final selection in the absence of the specific cut under consideration.
In figure 5, we present the distributions of the N-1 efficiencies plotted against the transverse momentum of the leading reconstructed muon ($p^{\mu}_{T}$) for the following conditions: $|E^{miss}-E^{\mu^{+}\mu^{-}}|/E^{\mu^{+}\mu^{-}}<0.4$ 5(a), $\Delta\phi_{\mu^{+}\mu^{-},\vec{E}^{\text{miss}}}>3.0$ 5(b), $\Delta R(\mu^{+}\mu^{-})<1.7$ 5(c), and $\text{cos}(\text{Angle}_{3D})<-0.8$ 5(d). These plots focus on the signal in the DH scenario (indicated by black closed circles), with $M_{Z^{\prime}}=30$ GeV, $\texttt{g}_{DM}=1.0$ and $\texttt{g}_{l}=0.003$, alongside standard model (SM) backgrounds marked with open colored markers.
The efficiency plots demonstrate that the application of these four stringent selection criteria effectively suppresses the background from the process $Z/\gamma\rightarrow\mu^{+}\mu^{-},\tau\tau$. This approach also minimizes contamination from events such as $WW$, $ZZ(4\mu)$, and $ZZ(2\mu 2\nu)$. Additionally, it ensures that the signal maintains a consistently flat efficiency across high transverse muon momentum ($p^{\mu}_{T}>40$ GeV).
5 Results
Figure 6: The recoil mass spectrum, for events passing the final selection listed in table 4, for the estimated SM backgrounds and different choices of the dark Higss masses generated based on the DH simplified model, with the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
Figure 6 shows the recoil mass spectrum for events that meet the final selection criteria listed in Table 4. The histograms represent the estimated SM backgrounds and various dark Higgs masses ($M_{h_{D}}=$ 30, 40, …, 80 GeV) generated based on the DH simplified model, with the coupling constants $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
The shape-based analysis utilizes the distributions of the recoil mass $(M_{rec})$, in different bins around the dark Higgs mass $(M_{h_{D}}-5<M_{rec}<M_{h_{D}}+5)$, for events passing the final event selection listed in table 4 as effective discriminators.
Table 5 presents the number of events that satisfy the final event selection criteria, calculated across different bins of the recoil mass $(M_{h_{D}}-5<M_{rec}<M_{h_{D}}+5)$ in line with the DH scenario. The assumed coupling constants are $\texttt{g}_{D}=1.0$ and $\texttt{g}_{l}=0.003$. The yields shown in the table encompass all anticipated relevant Standard Model backgrounds, based on an integrated luminosity of 10.8 ab-1 and a center-of-mass energy of $\sqrt{s}=240$ GeV. Furthermore, the table takes into account the total uncertainties, including both statistical and systematic components.
When analyzing an enriched sample of events for a specific signal process, it is essential to assess the entire distribution of a variable across all events rather than just counting the number of events within a specified signal region. In this framework, the profile likelihood can be employed for hypothesis testing by utilizing the asymptotic properties of the profiled likelihood ratio, defined as: $\lambda=\frac{L(b)}{L(s+b)}$. From this ratio, the p-value and statistical significance, denoted as $S_{L}$, can be derived using the following formula, SSL ; R2 ,
$S_{L}=\sqrt{-2ln\lambda},$
where $L(s+b)$ is the maximum likelihood value obtained in the full signal-plus-background unbinned maximum likelihood fit, and $L(b)$ is the maximum likelihood from the unbinned background-only fit.
The outcomes of this definition are presented in the last row of table 5, which quantifies the significance of the excess signal events in comparison to the standard model (SM) background events.
Figure 7: The statistical significance ($\it{S_{L}}$) versus integrated luminosity for $M_{h_{D}}=$ 30 (closed circles) and 60 (opened circles) GeV for events passing the final selection. The model signal represents the DH scenario with coupling constants $\texttt{g}_{l}=0.003$, and $\texttt{g}_{D}=1.0$ at $\sqrt{s}=240$ GeV. The dashed red line indicates $\it{S_{L}}=5$.
[…]
Table 5: The number of events satisfying the criteria of the final event selection are illustrated for several recoil mass bins $(M_{h_{D}}-5<M_{rec}<M_{h_{D}}+5)$, for each SM background, and the DH model signals for 10.8 ab-1 integrated luminosity and the coupling constants $g_{D}=1.0$, $g_{l}=0.003$. The total Uncertainty, including the statistical and systematic components, is indicated.
$M_{rec}$ (GeV) [15,25] [25,35] [35,45] [45,55] [55,65] [65,75] [75,85]
WWWW $179.1\pm 22.4$ $410.3\pm 45.8$ $648.6\pm 69.7$ $876.9\pm 92.6$ $1151.4\pm 120.0$ $1563.1\pm 161.2$ $1990.7\pm 204.0$
$ZZ\rightarrow 2\mu 2\nu$ $0.04\pm 0.2$ $0.32\pm 0.57$ $0.47\pm 0.69$ $1.51\pm 1.24$ $2.81\pm 1.70$ $5.18\pm 2.33$ $11.05\pm 3.50$
$ZZ\rightarrow 4\mu$ $04\pm 0.2 0.0$$14\pm 0.12 0.0$$41\pm 0.20 0.0$$27\pm 0.17 0.0$$41\pm 0.20 0.0$$55\pm 0.23 0.0$$041\pm 0.20$
Sum Bkgs $179.2\pm 22.4$ $410.6\pm 45.8$ $649.1\pm 69.7$ $878.4\pm 92.7$ $1154.2\pm 120.3$ $1568.3\pm 161.8$ $2001.8\pm 205.1$
DH signal $128.4\pm 17.1$ $2077.0\pm 212.6$ $2119.1\pm 216.9$ $2039.2\pm 208.9$ $1854.8\pm 190.4$ $1561.4\pm 161.1$ $832.4\pm 88.1$
$S_{L}$ $3.4\sigma$ $13.8\sigma$ $12.3\sigma$ $9.6\sigma$ $7.5\sigma$ $4.9\sigma$ $4.2\sigma$
We used the profile likelihood method to analyze our results statistically and performed a statistical test based on the $M_{rec}$ distributions. We used the modified frequentist construction CLs R58 ; R59 , which is based on the asymptotic approximation R2 , to derive exclusion limits on the product of signal cross sections and the branching fraction Br($Z^{\prime}$ $\rightarrow\mu\mu$) at a 95% confidence level. In the likelihood, the systematic uncertainties are treated as nuisance parameters.
In the mono-Z model, the 95% upper limit on the cross-section times the branching ratio for the DH simplified scenario is shown in figure 8. The result is presented for the muonic decay of the Z and with coupling constant values of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, for an integrated luminosity of 10.8 ab-1. This limit shows that the dark Higgs mass range from 20 to 80 GeV can be excluded for the coupling constant values of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
Figure 8: 95% CL upper limits on the cross-section times the branching ratio (expected), as a function of the dark Higgs mass ($M_{h_{D}}$) based on the mono-Z model, with the muonic decay of the Z. The black line represents the DH scenario with coupling constant values of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
6 Summary
The Future Circular Collider operating in Electron-Positron collision mode (FCC-ee) is an outstanding facility for uncovering particles that extend beyond the Standard Model (BSM). It offers a distinct signature for pinpointing unknown particles, such as dark matter, extra neutral gauge bosons, and Kaluza-Klein excitations.
Our research investigated the recoil mass distributions with low-mass dimuon pairs within the mono-Z model framework ($M_{rec}<90$ GeV), using simulated Monte Carlo (MC) data samples from the FCC-ee. These MC samples were generated from electron-positron collisions with $\sqrt{s}=240$ GeV, including signal and Standard Model background events. This setup corresponds to what is anticipated for FCC-ee Run 1, which is expected to have an integrated luminosity of 10.8 ab-1.
In this study, we investigated the effects of a simplified model scenario known as the dark Higgs (DH), focusing on dark Higgs production associated with a $Z^{\prime}$ boson at the FCC-ee. We analyzed the results from the muonic decay mode of the $Z^{\prime}$ boson, with coupling constants fixed at $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$.
We implemented effective discrimination cuts that entirely suppressed the $Z/\gamma$ background and significantly reduced the diboson $WW$ and $ZZ$ contaminations. This approach enabled us to differentiate more effectively between signal events and the SM backgrounds. Consequently, we achieved a noteworthy decrease in the SM backgrounds while maintaining the signal strength by applying the appropriate cuts, which are outlined in Table 4 for the DH scenario.
[…]
Finally, if the dark Higgs signal is not observed at the FCC-ee, we set upper limits on the mass of dark Higgs ($M_{h_{D}}$) at the 95% confidence level for the charged muonic decay channel of $Z^{\prime}$. Limits have been established for the dark Higgs scenario with $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, which excludes the dark Higgs mass range from 20 to 80 GeV.
This study highlights the promising opportunities to directly investigate the existence of a dark Higgs particle within the framework of physics that extends beyond the Standard Model at the FCC-ee. This is especially pertinent for the mono-$Z^{\prime}$ model, which features coupling parameters of $\texttt{g}_{l}=0.003$ and $\texttt{g}_{D}=1.0$, along with dark Higgs masses below 90 GeV. Notably, this domain has remained unexplored by the LHC.
Previous $e^{+}e^{-}$ colliders, including the LEP working group, updated the limits for the mass of a Higgs boson that decays invisibly using data from approximately 189-209 pb-1. No evidence was found for invisibly decaying Higgs bosons produced with a Z boson decaying into hadrons, electrons, or muons. The 95% confidence level limits exclude Higgs boson masses of 60 - 112.1 GeV from hadronic analysis and 60 - 91.3 GeV from leptonic analysis. Our analysis lowers the limit to 20 GeV.
The findings related to the electron-philic decay scenario of $Z^{\prime}$ would closely align with the results presented in this study, provided that the reconstruction efficiencies and resolutions for both electrons and muons are similar.
Acknowledgements.
The author of this paper would like to thank Tongyan Lin, co-author of R1 , for providing us with the UFO model files, helping us to generate the signal events, and cross-checking the results. In addition, this paper is based on works supported by the Science, Technology, and Innovation Funding Authority (STDF) under grant number 48289.